# Maximal information at the edge of stability in excitatory-inhibitory neural populations

### v1

@@ Abstract
where $p_{h}^{\mathrm{st}}$ is the distribution of the strength of the stimulus, and $H_{\bm{x}|h}$ is the conditional differential entropy of $\bm{x}$ at a given input value $h$. We assume that $p_{h}^{\mathrm{st}}$ is fully characterized by the mean input strength, $\mu_{h}$, and its variance, $\sigma_{h}$, so that $p_{h}=\mathcal{N}(\mu_{h},\sigma_{h})$. Then, we have that
$I_{\bm{x},h}(t)$ $=\dfrac{1}{2}\log\dfrac{\det\left[\Sigma^{\mathrm{st}}+K(t)\begin{pmatrix}\sigma_{h}^{2}&0\\ 0&0\end{pmatrix}K(t)^{T}\right]}{\det\left(\Sigma^{\mathrm{st}}\right)}$ (8)
@@
where $\Sigma^{\mathrm{st}}$ is the covariance matrix of the unperturbed system at steady state, and $K(t)$ is the time-dependent gain matrix that we derive in the Supplemental Material suppinfo_ref . In Figure 3a, we plot the time evolution of the mutual information. At long times, we find once more that the mutual information is maximized at the edge of stability (Figure 3b), with $I_{\bm{x},h}$ diverging as $k\to k_{c}$. We note that, since the differential entropy for the continuous input distribution is not necessarily positive, the bounds in Eq. (5) cannot be straightforwardly applied. In particular, while the maximal information content of the input was associated with its switching dynamics in the previous setting, there is now no a priori limit to the information that the system can encode.
@@
The scenario becomes more intricate at short times after the stimulus. In the inhibition-stabilized regime, where $w>1$, the response of the neural populations exhibits a faster increase for stronger excitatory couplings away from the edge of instability. To assess the system’s responsiveness, we introduce a metric of sensitivity defined as $\chi_{\bm{x},h}=\partial_{t}^{2}I_{\bm{x},h}(t)|_{t=t_{\mathrm{stim}}}$, which quantifies the instantaneous increase in mutual information when the stimulus arrives at time $t=t_{\mathrm{stim}}$ derivative_note . In Figure 3c, we show that $\chi_{\bm{x},h}$ peaks at an optimal inhibition strength $k_{\rm max}>1$, whose complete expression is given in suppinfo_ref . Crucially, this optimal value depends on the excitation strength as well, i.e., $k_{\rm max}\equiv k_{\rm max}(w)$ (see Fig. 3d). This highlights that the inhibition regime for the optimal response at short times is drastically different from that at long times (see Figs. 3d-e). In particular, since $k_{c}(w)<k_{\mathrm{max}}(w)$ for all $w$, our results unravel a fundamental trade-off between achieving maximum accuracy and the speed at which the neural populations encode information about the external stimulus, akin to a speed-accuracy trade-off lan2012energy . As we show in Figure 3f, long-time accuracy is generally achieved at lower inhibition, whereas sensitivity maximization requires a larger value of $k$.
@@
Overall, our analysis marks a significant step towards the understanding of how the structure of connectivity shapes information encoding in neuronal population dynamics. We showed analytically in an exactly solvable regime that the mutual information between an input and the receiving neuronal populations is maximized at the edge of stability, highlighting the importance of the balance between excitation and inhibition. As non-normal synaptic interactions are crucial for achieving this optimal state, our findings confirm that the underlying structural connectivity is essential for the emergence of critical behaviors that support optimal information encoding and complex dynamical evolution both in whole-brain connectomes barzon2022criticality and in artificial recurrent neural networks kerg2019non . Notably, alterations in excitatory-inhibitory balance have been experimentally related to the loss of information-processing efficiency observed in pathological conditions dehghani2016dynamic ; sohal2019excitation . Further, our results point towards the hypothesis that the bran might be poised at criticality to optimize its information-processing capabilities beggs2008criticality ; o2022critical ; mariani2021b ; mariani2022disentangling , and that computation at edge-of-chaos may be optimal langton1990computation .
@@
Our predictions are also consistent with recent experimental studies in which theoretical tools from response theory have been applied to extensive whole-brain neuronal recordings. The emergent dynamics of several brain regions has been shown to lie at the edge of stability dahmen2019second ; morales2023quasiuniversal , with a distance from instability that only slightly varied along the cortex. Such heterogeneity might be explained as an increase in the inhibition level wang2020macroscopic . Our findings suggest that this observed feature may be related to the tuning of sensitivity to different timescales murray2014hierarchy ; manea2022intrinsic . Importantly, the external stimulus considered here may be immediately generalized to a high-dimensional signal representing, for example, multiple stimuli with different characteristics (e.g., frequency, intensity) targeting spatially separated populations.
@@
Although we focused on a paradigmatic - yet widely used - model, our approach applies in principle to any neuronal connectivities and can be used to investigate more detailed synaptic structures and their effects on information-processing performances. Although it will be crucial to understand how strongly non-linear regimes affect the behavior of the mutual information and the information-theoretic features of the system in general, our work paves the way to the unraveling of the fundamental mechanisms supporting information encoding and sensitivity in neuronal networks.
@@ Acknowledgements.
G.N. acknowledges funding provided by the Swiss National Science Foundation through its Grant CRSII5_186422.

### v2

@@ Abstract
where $p_{h}^{\mathrm{st}}$ is the distribution of the strength of the stimulus, and $H_{\bm{x}|h}$ is the conditional differential entropy of $\bm{x}$ at a given input value $h$. We assume that $p_{h}^{\mathrm{st}}$ is fully characterized by the mean input strength, $\mu_{h}$, and its variance, $\sigma_{h}$, so that $p_{h}=\mathcal{N}(\mu_{h},\sigma_{h})$. Then, we have:
$I_{\bm{x},h}(t)=\dfrac{1}{2}\log\dfrac{\det\left[\hat{\Sigma}^{\mathrm{st}}+\hat{K}(t)\begin{pmatrix}\sigma_{h}^{2}&0\\ 0&0\end{pmatrix}\hat{K}(t)^{T}\right]}{\det\left(\hat{\Sigma}^{\mathrm{st}}\right)}$ (8)
@@
where $\hat{K}(t)$ is the time-dependent gain matrix that we derive in the SM suppinfo_ref . In Fig. 3a, we plot the time evolution of the mutual information. At long times, we find once more that the mutual information is maximized at the edge of stability (Fig. 3b), with $I_{\bm{x},h}$ diverging as $k\to k_{c}$. We note that, since the differential entropy for the continuous input distribution is not necessarily positive, the bounds in Eq. (5) cannot be straightforwardly applied. In particular, while the maximal information content of the input was associated with its switching dynamics in the previous setting, there is now no a priori limit to the information that the system can encode.
The scenario becomes more intricate at short times after the stimulus. In the inhibition-stabilized regime, where $w>1$, the response of the neural populations exhibits a faster increase for stronger excitatory couplings away from the edge of instability. To assess the system’s responsiveness, we introduce a metric of sensitivity:
$\chi_{\bm{x},h}$ $=\frac{\partial^{2}I_{\bm{x},h}(t)}{\partial t^{2}}\biggr{|}_{t=t_{\mathrm{stim}}}$ (9)
$=\frac{\sigma_{h}^{2}}{2D}\frac{[r+w(k-k_{c})][2wk_{c}^{2}+r(k+1)]}{2r(k-k_{c})+w(1+k^{2})}$
@@
which measures how quickly the information on the stimulus increases immediately after the stimulation time $t=t_{\mathrm{stim}}$ derivative_note . In Fig. 3c, we show that $\chi_{\bm{x},h}$ peaks at an optimal inhibition strength $k_{\rm max}(w)>1$, whose complete expression is given in the SM suppinfo_ref . Crucially, this optimal value depends on the excitation strength (see Fig. 3d). This reveals that the inhibition regime for the optimal response at short times is drastically different from that at long times, as we show in Figs. 3d-e. In particular, since $k_{c}(w)<k_{\mathrm{max}}(w)$ for all $w$, our results unravel a fundamental trade-off between achieving maximum accuracy and the speed at which the neural populations encode information about the external stimulus, akin to a speed-accuracy trade-off emerging in different biological contexts lan2012energy . Remarkably, similar trade-offs have been recently found by studying the response of recurrent neural networks with random connectivities at different observation times azizpour2023available . As we show in Fig. 3f, long-time accuracy is generally achieved at lower inhibition, whereas sensitivity maximization requires a larger value of $k$. In the SM suppinfo_ref , we show that, when both populations receive the prolonged stimulus, sensitivity is maximized at $k\to\infty$, suggesting that inhibition-dominated networks exhibit an enhanced tracking ability of the input renart2010asynchronous .
@@
Overall, our analysis marks a significant step towards the understanding of how the structure of connectivity shapes information encoding in neuronal population dynamics. We have shown, both analytically in an exactly solvable regime and numerically for a nonlinear scenario, that the mutual information between a switching input and the receiving neuronal populations is controlled by the balance between the excitatory and inhibitory couplings, and peaks at the edge of linear stability. Moreover, we found that an increased inhibition strength is instrumental for establishing a robust response at short times, highlighting the importance of precisely tuning excitation and inhibition to achieve optimal encoding at different timescales. As non-normal synaptic interactions are crucial for realizing this optimal state, our findings at a coarse-grained level underscore the essential role of the underlying connectivity. In particular, structural connectivity has been shown to be essential in supporting complex dynamical evolution both in whole-brain connectomes barzon2022criticality and in artificial recurrent neural networks baggio2020efficient . Our study opens the avenue for an information-theoretic quantification of these emerging features from first principles, at the level of neural populations. In future works, the extension of similar ideas to more microscopic models might shed light on the intimate link between complex, even chaotic langton1990computation ; clark2024theory , dynamics, and information processing performances.
@@
Notably, alterations in excitatory-inhibitory balance have been experimentally related to the loss of information-processing efficiency observed in pathological conditions dehghani2016dynamic ; sohal2019excitation . Our predictions are also consistent with recent experimental studies in which theoretical tools from response theory have been applied to extensive whole-brain neuronal recordings. The emergent dynamics of several brain regions has been shown to lie at the edge of stability dahmen2019second ; morales2023quasiuniversal , with a distance from instability that only slightly varied along the cortex. Such heterogeneity might be explained as an increase in the inhibition level wang2020macroscopic , and our findings suggest that this observed feature may be related to the tuning of sensitivity to different timescales murray2014hierarchy ; manea2022intrinsic . Importantly, the external input considered here may be immediately generalized to a high-dimensional signal representing, for example, multiple stimuli with different characteristics (e.g., frequency, intensity) targeting spatially separated populations potentially evolving on different timescales.
@@
Although we focused on a paradigmatic - yet widely used - model, our approach can be extended to investigate more detailed and microscopic synaptic structures and include the role played by plasticity to drive accurate encoding. Overall, our work paves the way to the unraveling of the fundamental mechanisms supporting information encoding and sensitivity in neuronal networks.
@@ Acknowledgements.
Acknowledgments.—G.N. acknowledges funding provided by the Swiss National Science Foundation through its Grant CRSII5_186422.